The displacement vectors A and B, when added together, give the resultant vector R, so that R = A + B. Use the data in the drawing and the fact that φ = 31° to find the magnitude R of the resultant vector and the angle θ that it makes with the +x axis.


HOW CAN I SOLVE THIS?



1
Expert's answer
2022-01-25T06:50:14-0500

To find the magnitude R of the resultant vector, we need three things: the magnitude "|\vec A|" of A, the magnitude "|\vec B|" of B, and the angle "\phi" between them:


"A_x=|\vec A|,\\
B_x=|\vec B|\cos\phi,\\
A_y=0,\\
B_y=|\vec B|\sin\phi. \\\space\\
R_x=A_x+B_x=|\vec A|+|\vec B|\cos\phi,\\
R_y=A_y+B_y=|\vec B|\sin\phi,\\\space\\
|\vec R|=\sqrt{(R_x)^2+(R_y)^2}."


To find the angle "\theta" of R above the x-axis:


"\theta=\arctan\dfrac{R_y}{R_x}=\arctan\dfrac{|\vec B|\sin\phi,}{|\vec A|+|\vec B|\cos\phi}."